OpenStudy (anonymous):

MEDAL Simplify the expression the fraction with numerator of the square root of negative twenty five and denominator of the quantity five minus two times i plus the quantity one minus three times i

OpenStudy (hybrik):

\[-\sqrt{25}/5-2(i)+1-3(i)\]

OpenStudy (anonymous):

that is the denominator

OpenStudy (hybrik):

You would get \[\frac{ -5 }{ 5-2i+1-3i}\]

OpenStudy (hybrik):

\[\frac{ -5 }{ 5+1-2i-31 } = \frac{ -5 }{ 6-5i }\]

OpenStudy (hybrik):

The answer is \[\frac{ -5 }{ 6-5i }\]

OpenStudy (anonymous):

that is not an answer choice

OpenStudy (hybrik):

Can you show me the choices?

OpenStudy (hybrik):

To be honest I messed up on the numerator, its not -sqr25

OpenStudy (hybrik):

You cant simplify \[\sqrt{-25}\]

OpenStudy (anonymous):

i guess the answer is 5+20i/17

OpenStudy (hybrik):

Can you take a screenshot of the normal, or is it actually written like that?

OpenStudy (hybrik):

Choices?

OpenStudy (anonymous):

Simplify the expression tfraction with numerator of the square root of negative twenty five and denominator of the quantity five minus two times i plus the quantity one minus three times i. the quantity five plus twenty times i divided by seventeen the quantity negative twenty five plus thirty times i divided by sixty the quantity twenty five plus thirty times i divided by sixty one the quantity negative five plus twenty times i divided by seventeen